Please answer all questions.

1. One simple model for a person running the 100 m dash is to assume the sprinter runs with constant acceleration until reaching top speed, then maintains that speed through the finish line. If a sprinter reaches his top speed of 11.4 m/s in 2.34 s , what will be his total time?

2. Excellent human jumpers can leap straight up to a height of 120 cm off the ground. To reach this height, with what speed would a person need to leave the ground? Express your answer to two significant figures and include the appropriate units.

3. A basketball player can jump to a height of 1.1 m off the ground, so he can easily dunk a basketball. For such a leap, what is his "hang time" - the time spent in the air after leaving the ground and before touching down again? Express your answer to two significant figures and include the appropriate units.

4. When striking, the pike, a predatory fish, can accelerate from rest to a speed of 3.8 m/s in 0.14 s . A) What is the acceleration of the pike during this strike? Express your answer using two significant figures. B) How far does the pike move during this strike? Express your answer using two significant figures.
Answer #1

1) By using,

v = u + at

a = v/t = 11.4/2.34 = 4.87 m/s^2

s = ut + (1/2)at^2

= 0 + (1/2)*4.87*(2.34)^2

= 13.34 m

Now, t' = (100 - 13.34)/11.4 s = 7.6 s

Total time = 7.6 + 2.34

= 9.94 s